OpenStudy (anonymous):

Pablo is making ramps of different lengths. The ramps are made with rectangular and square wooden boards. The relationship between the length of the ramp in feet, f(n), and the number of square boards used, n, is represented by the function below. f(n) = 20n + 3 , where 2 ≤ n ≤ 6 Which statement is correct about the range of the length of the ramp for the given domain? It varies from 2 to 6. It varies from 43 to 123. It varies from 40 to 120. It varies from 2 to 123.

OpenStudy (anonymous):

It varies from 43 to 123.

OpenStudy (anonymous):

thanks (:

OpenStudy (anonymous):

You're welcome. Besides if you want an explanation just ask for it.

OpenStudy (anonymous):

yes please, I would like to know how to do it also, instead of just the answer (:

OpenStudy (anonymous):

You know that n can be 2, 3, 4, 5 or 6. The domain is basicly the maximum and minimum of square boards used. Thus replace n by 2 and by 6: f(2) = 20 * 2 + 3 = 43 f(6) = 20 * 6 + 3 = 123

OpenStudy (anonymous):

Here's the mathematical notation of the domain [43;123]

OpenStudy (anonymous):

thanks so much :D your awesome.